OpenStudy (anonymous):

Please help If the quadratic formula is used to solve 2x(x + 5) = 4, what are the solutions?

OpenStudy (anonymous):

2x(x + 5) = 4 2x^2+10x=4 2x^2+10x-4=0

OpenStudy (anonymous):

Use Quadracic formula to solve. Do you know the formula?

OpenStudy (anonymous):

Yes I do but this one just gets me @Linchpin

OpenStudy (anonymous):

Okay,Can you draw the formula here?

OpenStudy (anonymous):

Or should I?

OpenStudy (anonymous):

Yup